Consider the reaction of calcium hydroxide and hydrochloric acid. If 63.48 g of the calcium based product forms from 44.50 g of hydrochloric acid, what is the percentage yield?


1
Expert's answer
2021-03-10T17:17:44-0500

Ca(OH)2 + 2HCl = CaCl2 + 2H2O

m(CaCl2) real = 63.48 g

m(CaCl2) = 44.50 / 36.45 / 2 × 110.98 = 67.74 g

Yield = 63.48/67.74 × 100% = 93.71%